OpenStudy (anonymous):

26.While shopping you found a pair of pants on sale for 1/4 off the regular price. If the regular price is $35.50, how much will be marked off the price? Round to the nearest cent. A.$8.88 B.$8.00 C.$8.45 27.True or False: (1/3)(3n + 3) = n + 1

OpenStudy (anonymous):

Medal if you help.

OpenStudy (anonymous):

26. $35.50 * 0.25 = $8.88 27. (1/3)(3n+3) = n + 1 = true

OpenStudy (anonymous):

26 8.88 27. true